ABCD is parallelogram.AE and BE are angle bisector angle A and angle B . show that angle AEB is 90 degree
Answer by mananth(16946)

AE bisects angle A
let each bisected angle be x
angle A = 2x
Similarly let y be each of the bisected angle
2y = angle B
2x+2y = 180 deg ( adjacent angles of a parallelogram)
x+y = 90
In triangle ABE
x +y + angle AEB = 180 deg. ( sum of angles of a triangle)
90+angle AEB = 180 deg
angle AEB = 180-90
Angle AEB = 90 deg